A child starts sledding from rest, going down a 40-m-long 7.5° incline, then coasting across a horizontal stretch. The mass of sled + child is 35 kg, and the coefficient of kinetic friction is 0.060.
What is the total time for the ride?
1
Expert's answer
2020-12-14T07:24:59-0500

"\text {inclined track:}"

"F_1=ma_1"

"a_1=g(\sin \theta -\mu \cos \theta )"

"a_1 = 9.8(sin7.5^0-0.06*cos 7.5^0) = 0.7"

"s = V_0t_1+\frac{a_1t_1^2}{2};V_0=0"

"t_1 = \sqrt{\frac{2s}{a_1}}=\sqrt{\frac{2*40}{0.7}}=10.7"

"V_1 = V_0+at_1"

"V_1=0.7*10.7 =7.49"


"\text{horizontal track:}"

"F_2=ma_2"

"a_2= \mu*g"

"V_2= V_0-at_2"

"V_2=0 \text{ stop motion}"

"V_0 = 7.49 \text{ speed at the end inclined track}"

"t_2= \frac{V_0}{a_2}=\frac{V_0}{\mu*g}=\frac{7.49}{9.8*0.06}=12.7"


"\text{All time:}"

"t = t_1+t_2 = 12.7+10.7 = 23.4"

Answer:total time 23.4 seconds